The life length of a certain brand A of an electronic device is normally distributed with mean
40 and standard deviation 3, while that of another brand B is normally distributed with mean
45 and standard deviation 6. Which of the two brands should be preferred for a continuous use
of
(a) 45 hours ?
(b) 55 hours ?
1
Expert's answer
2021-08-15T18:11:42-0400

Let "X=" the life length of a certain brand A of an electronic device: "X\sim N(\mu_1, \sigma_1^2)."

Let "Y=" the life length of a certain brand B of an electronic device: "Y\sim N(\mu_2, \sigma_2^2)."

Given


"\mu_1=40\ h, \sigma_1=3\ h,"

"\mu_2=45\ h, \sigma_2=6\ h"

(a)


"P(X\geq45)=1-P(X<45)"

"=1-P(Z<\dfrac{45-40}{3})\approx1-P(Z<1.6667)"

"\approx0.0478"


"P(Y\geq45)=1-P(Y<45)"

"=1-P(Z<\dfrac{45-45}{6})=1-P(Z<0)"

"=0.5"

Brands B should be preferred for a continuous use of 45 hours.


(b)


"P(X\geq55)=1-P(X<55)"

"=1-P(Z<\dfrac{55-40}{3})\approx1-P(Z<5)"

"\approx0.0000003"




"P(Y\geq55)=1-P(Y<55)"

"=1-P(Z<\dfrac{55-45}{6})=1-P(Z<1.6667)"

"\approx0.0478"

Brands B should be preferred for a continuous use of 55 hours.